# Scalable Multi-Task Learning for Particle Collision Event Reconstruction with Heterogeneous Graph Neural Networks

### v2

@@ 2 Related work
More recently, several efforts have been made towards a fully inclusive reconstruction of beauty-hadron decays with GNNs at LHCb and Belle II [26, 27, 22]. These developments were prompted by Kahn et al. [26] with the introduction of a novel edge classification target for hierarchical decay chains known as the lowest common ancestor generations (LCAG) matrix. This compact representation enables one to learn the hierarchical structure of a decay solely from its final-state particles. For each edge relation between final-state particles, a multi-class label is used, which signifies the generational class of the shared ancestor.
@@ 3.1 Graph Neural Networks
{split}$\split e^{\prime k}=\phi^{e}(e^{k},v^{r_{k}},v^{s_{k}},u)\qquad&\bar{e}^{\prime i}=\rho^{e\rightarrow v}(\{E^{\prime i}\})\\ v^{\prime i}=\phi^{v}(\bar{e}^{\prime i},v^{i},u)\qquad\bar{e}^{\prime}=\rho^{e\rightarrow u}(\{E^{\prime}\})\\ u^{\prime}=\phi^{u}(\bar{e}^{\prime},\bar{v}^{\prime},u)\qquad\bar{v}^{\prime}=\rho^{v\rightarrow u}(\{V^{\prime}\})\$ (1)
The edge update function $\phi^{e}$ is a learnable multi-layer-perceptron (MLP), that takes as input the existing edge representation $e^{k}$ for edge $k$, adjoining receiver and sender node representations ($v^{r_{k}}$ and $v^{s_{k}}$) and a global representation $u$. The subsequent node update with MLP $\phi^{v}$ for node $i$, takes as input the node $v^{i}$ representation, the edge-to-node aggregation $\bar{e}^{\prime i}$ and $u$. Here, $\bar{e}^{\prime i}$ aggregates edge representations for the set of edges $E^{\prime i}$ that node $i$ receives using an aggregation function $\rho^{e\rightarrow v}$. Finally, sets of edges and nodes, $E^{\prime}$ and $V^{\prime}$, are aggregated globally to give $\bar{e}^{\prime}$ and $\bar{v}^{\prime}$ using the aggregation functions $\rho^{e\rightarrow u}$ and $\rho^{v\rightarrow u}$. The global update $u^{\prime}$ with MLP $\phi^{u}$ takes as input global aggregations of edges ($\bar{e}^{\prime}$) and nodes ($\bar{v}^{\prime}$) and the existing global representation, $u$.
@@ 3.2 HGNN layer for particle reconstruction and pruning
Figure 1: (a) Heterogeneous graph representation (above) and HGNN architecture (below) for a simultaneous beauty hadron reconstruction and PV association. In the graph solid lines indicate all true edges and their class labels, while dotted lines provide examples of false edges with label zero. (b) Heterogeneous modification in blue to the GNN layer updates from Battaglia et al[45] in red. The addition of pruning tasks is shown in light brown.
@@
{split}$\split e^{\prime k}_{\rm tr}=\phi^{e_{\rm tr}}(e^{k}_{\rm tr},v^{r_{k}}_{\rm tr},v^{s_{k}}_{\rm tr},u)\qquad&\bar{e}^{\prime i}_{\rm tr}=\rho^{e_{\rm tr}\rightarrow v_{\rm tr}}(\{E^{\prime i}_{\rm tr}\})\\ e^{\prime l}_{\rm pv-tr}=\phi^{e_{\rm pv-tr}}(e^{l}_{\rm pv-tr},v^{m_{l}}_{\rm tr},v^{n_{l}}_{\rm pv},u)\qquad\bar{e}^{\prime i}_{\rm pv-tr}=\rho^{e_{\rm pv-tr}\rightarrow v_{\rm tr}}(\{E^{\prime i}_{\rm pv-tr}\})\\ v^{\prime i}_{\rm tr}=\phi^{v_{\rm tr}}(v^{i}_{\rm tr},\bar{e}^{\prime i}_{\rm tr},\bar{e}^{\prime i}_{\rm pv-tr},u)\qquad\bar{e}^{\prime j}_{\rm pv-tr}=\rho^{e_{\rm pv-tr}\rightarrow v_{\rm pv}}(\{E^{\prime j}_{\rm pv-tr}\})\\ v^{\prime j}_{\rm pv}=\phi^{v_{\rm pv}}(v^{j}_{\rm pv},\bar{e}^{\prime j}_{\rm pv-tr},u)\qquad\bar{e}^{\prime}_{\rm tr}=\rho^{e_{\rm tr}\rightarrow u}(\{E^{\prime}_{\rm tr}\}),\bar{v}^{\prime}_{\rm tr}=\rho^{v_{\rm tr}\rightarrow u}(\{V^{\prime}_{\rm tr}\})\\ u^{\prime}=\phi^{u}(\bar{e}^{\prime}_{\rm tr},\bar{e}^{\prime}_{\rm pv-tr},\bar{v}^{\prime}_{\rm tr},\bar{v}^{\prime}_{\rm pv},u)\qquad\bar{e}^{\prime}_{\rm pv-tr}=\rho^{e_{\rm pv-tr}\rightarrow u}(\{E^{\prime}_{\rm pv-tr}\}),\bar{v}^{\prime}_{\rm pv}=\rho^{v_{\rm pv}\rightarrow u}(\{V^{\prime}_{\rm pv}\})\;,$ (2)
@@
Pruning can be performed either via a top-$k$ operation or by introducing a threshold $y_{\rm cut}$. In this paper, we adopt the latter approach, given its better observed performance and greater interpretability in terms of purity. Furthermore, we focus on the application of pruning during inference to boost the speed and remove redundant background edges and nodes to enhance reconstruction performance. We acknowledge that integrating pruning into training (i.e., ‘pruning-aware’ training, a form of dynamic graph construction) could further improve the reconstruction. Despite not employing pruning-aware training, our empirical results show that incorporating extra MLPs per node and edge improves the performance of the algorithm, as detailed in the subsequent sections.
@@ 3.3 HGNN architecture and objective loss
{split}=&$\split\mathcal{L}=&\mathcal{L}_{\rm CE}\left(e_{\rm out},y^{\rm LCA}\right)+\beta^{e_{\rm tr}}\sum^{n^{\rm GNN}_{\rm layers}}_{i}\mathcal{L}_{\rm BCE}\left(\hat{y}^{e_{\rm tr}}_{i},y^{e_{\rm tr}}_{i}\right)+\beta^{v_{\rm tr}}\sum^{n^{\rm GNN}_{\rm layers}}_{i}\mathcal{L}_{\rm BCE}\left(\hat{y}^{v_{\rm tr}}_{i},y^{v_{\rm tr}}_{i}\right)\\ +\beta^{e_{\rm pv-tr}}\sum^{n^{\rm GNN}_{\rm layers}}_{i}\mathcal{L}_{\rm BCE}\left(\hat{y}^{e_{\rm pv-tr}}_{i},y^{e_{\rm pv-tr}}_{i}\right)$ (3)
@@ 3.4 Datasets and training procedure
As depicted in Figure 1, collision events are represented as a heterogeneous graph. Each node and edge type has an input-feature representation. For charged particles (tracks), the input features include the track origin point, $(x_{\rm tr},y_{\rm tr},z_{\rm tr})$, track momenta ($p_{x},p_{y},p_{z}$) and track charge $q$. Meanwhile, PV nodes are represented by their position coordinates, $(x_{\rm pv},y_{\rm pv},z_{\rm pv})$. In contrast, for the GNN benchmarks, we adopted DFEI’s approach of appending the coordinates of the PV with the minimum impact parameter to the features of each track. Although highly discriminating timing information is anticipated for PV association in LHCb Upgrade II, this study limited its scope to using only the previously defined positional and kinematic features.
@@
Initially, a fully connected graph is defined between all track and PV nodes with track-track and PV-track edges. The input edge representation for PV-track edges includes only the impact parameter of the track with respect to the PV. Meanwhile, track-track edge features include the angle $\theta$ between the three-momentum directions of the two particles, the difference $\Delta z_{\rm tr}$ in origin $z_{\rm tr}$, a Boolean indicating a shared PV according to the minimum impact parameter, and the momentum-transverse distance, which is the distance between the origin points in a plane transverse to the momentum direction. We adopted a loose prefiltering of track-track edges with 99% efficiency from the DFEI [22], which requires edges to satisfy $\theta<0.26$ rad. or the condition of a shared PV according to the minimum impact parameter. For the remaining edges, the ground-truth LCAG is determined $y^{\rm LCA}$, where $y^{\rm LCA}=0$ indicates no shared ancestors. The non-zero values were restricted to $y^{\rm LCA}=1,2,3$, which were found to be sufficient to represent the decay hierarchies present in the simulation dataset. Finally, global input features include the multiplicity of tracks and PVs in the event.
@@
The training and validation datasets consisted of 40,000 and 10,000 events respectively, which are equivalent to those used in our previous work. Nominal training for the ablation studies was performed on an NVIDIA L40S GPU with a batch size of 12. For ablation studies, the base schedule consisted of 30 epochs at a learning rate of $\alpha=10^{-3}$, followed by two epochs at $\alpha=10^{-4}$. Before evaluating the reconstruction performance of the selected architectures, we fine-tuned further with two additional epochs at $\alpha=5\times 10^{-5}$ and two epochs at $\alpha=2\times 10^{-5}$. During all training phases, training and validation losses as well as classification accuracies were monitored (see  A for examples of training and validation loss curves). The final performance was measured using a test set of 10,000 independent inclusive beauty-hadron decay events and multiple exclusive decay datasets, each containing 5,000 events.
@@ 4.1 Ablation studies
4.2 Pruning performance
Figure 2(a) and (b) show the discrimination power of the track edge and node scores $\hat{y}^{e_{\rm tr}}$ and $\hat{y}^{v_{\rm tr}}$ at HGNN layers 1, 2, 3 and 8. The discrimination power increases significantly between layers 1 and 2 particularly for the edges, with the area under the curve (AUC) increasing from 0.973 to 0.999. While the discrimination power saturates with subsequent layers, there are marginal increases in the discrimination power for each subsequent HGNN layer.
@@ 4.2 Pruning performance
Figure 2: (a) and (b) Receiver Operating Characteristic curves for edge and node scores $\hat{y}^{e_{\rm tr}}$ and $\hat{y}^{v_{\rm tr}}$ at layers 1, 2, 3 and 8. The pruning selections used in the subsequent results are highlighted. (c) Relationship between the number of particles from a beauty hadron and the number of particles selected with a tight edge and node pruning $\hat{y}^{e_{\rm tr}}_{8}>0.2$ and $\hat{y}^{v_{\rm tr}}_{8}>0.2$. (d) Signal efficiency and background retention as functions of track multiplicity for tight and loose ($\hat{y}^{e_{\rm tr}}_{8}>0.01$ and $\hat{y}^{v_{\rm tr}}_{8}>0.01$) pruning selections.
Figure 3: GNN and HGNN GPU (left) and CPU (right) mean inference time as a function of track multiplicity per event with and without early pruning ($\hat{y}^{e_{\rm tr}/v_{\rm tr}}_{1-3}>10^{-4}$). The error bands indicate ±$\pm 1\sigma$, where $\sigma$ is the standard deviation of the inference time.
Figure 2(c) and (d) further quantify the performance of pruning after applying a selection for the LCAG class pediction to be above zero. In particular, Figure 2(c) displays the relationship between the number of tracks from beauty hadrons and the number of tracks selected with a tight pruning selection, $\hat{y}^{e_{\rm tr}}_{8}>0.2$ and $\hat{y}^{v_{\rm tr}}_{8}>0.2$. Meanwhile, Figure 2(d) compares the efficiencies for signal and background tracks, as a function of track multiplicity using the tight pruning selection and a looser pruning selection with $\hat{y}^{e_{\rm tr}}_{8}>0.01$ and $\hat{y}^{v_{\rm tr}}_{8}>0.01$. The efficiencies are defined by the ratio of surviving signal and background tracks after pruning and LCAG selections.
The node and edge pruning layers aim to accelerate the inference time of the models, particularly the scaling of inference time with track multiplicity. Figure 3 shows GPU and single-threaded CPU inference times, measured with a NVIDIA RTX 4090 GPU and a Intel Core i9-14900K 3.2 GHz CPU, as a function of track multiplicity for the cases of no pruning selection and a loose edge pruning selection at layers 1-3 ($\hat{y}^{e_{\rm tr}/v_{\rm tr}}_{1-3}>10^{-4}$). Early edge pruning yields noticeable speed-ups above approximately 250 tracks; at multiplicities exceeding 400, we observe a 2-3$\times$ acceleration on GPU and a 5$\times$ acceleration on CPU. For example, for events with track multiplicity around 300, single-threaded CPU inference with early pruning takes approximately 300 ms compared to the 2.2 s timing reported for the multi-stage GNN approach in DFEI [22] on a 2.2 GHz Intel Core processor.

### v3

@@ 2 Related work
More recently, several efforts have been made towards a fully inclusive reconstruction of beauty-hadron decays with GNNs at LHCb and Belle II [26, 27, 22]. These developments were prompted by Kahn et al. [26] with the introduction of a novel edge classification target for hierarchical decay chains known as the lowest common ancestor generations (LCAG) matrix. This compact representation enables one to learn the hierarchical structure of a decay solely from its final-state particles. For each edge relation between final-state particles, a multi-class label is used, which signifies the generational class of the shared ancestor. Kahn et al. further benchmarked their GNN-based approach against transformer architectures for LCAG reconstruction, demonstrating significantly better performance of GNNs for this hierarchical task.
@@ 3.1 Graph Neural Networks
$split e^{\prime k}=\phi^{e}(e^{k},v^{r_{k}},v^{s_{k}},u)\qquad&\bar{e}^{\prime i}=\rho^{e\rightarrow v}(\{E^{\prime i}\})\\ v^{\prime i}=\phi^{v}(\bar{e}^{\prime i},v^{i},u)\qquad&\bar{e}^{\prime}=\rho^{e\rightarrow u}(\{E^{\prime}\})\\ u^{\prime}=\phi^{u}(\bar{e}^{\prime},\bar{v}^{\prime},u)\qquad&\bar{v}^{\prime}=\rho^{v\rightarrow u}(\{V^{\prime}\})\\$ (1)
The edge update function $\phi^{e}$ is a learnable multi-layer-perceptron (MLP), that takes as input the existing edge representation $e^{k}$ for edge $k$, adjoining receiver and sender node representations ($v^{r_{k}}$ and $v^{s_{k}}$) and a global representation $u$. The subsequent node update with MLP $\phi^{v}$ for node $i$, takes as input the node $v^{i}$ representation, the edge-to-node aggregation $\bar{e}^{\prime i}$ and $u$. Here, $\bar{e}^{\prime i}$ aggregates edge representations for the set of edges $E^{\prime i}$ that node $i$ receives using an aggregation function $\rho^{e\rightarrow v}$  [^1] For example, if node $1$ receives edges from nodes $2$ and $3$, the incoming edges $(2\!\to\!1)$ and $(3\!\to\!1)$ are first updated as $e^{\prime 21}=\phi^{e}(e^{21},v^{1},v^{2},u)$ and $e^{\prime 31}=\phi^{e}(e^{31},v^{1},v^{3},u)$. These are then aggregated, for instance $\bar{e}^{\prime 1}=e^{\prime 21}+e^{\prime 31}$ for sum pooling, and the node representation is updated as $v^{\prime 1}=\phi^{v}(v^{1},\bar{e}^{\prime\,1},u)$.. Possible aggregation functions include non-parametric functions such as sum, mean, and max pooling. Finally, sets of edges and nodes, $E^{\prime}$ and $V^{\prime}$, are aggregated globally to give $\bar{e}^{\prime}$ and $\bar{v}^{\prime}$ using the aggregation functions $\rho^{e\rightarrow u}$ and $\rho^{v\rightarrow u}$. The global update $u^{\prime}$ with MLP $\phi^{u}$ takes as input global aggregations of edges ($\bar{e}^{\prime}$) and nodes ($\bar{v}^{\prime}$) and the existing global representation, $u$.
@@ 3.2 HGNN layer for particle reconstruction and pruning
Figure 1: (a) Heterogeneous graph representation (above) and HGNN architecture (below) for a simultaneous beauty hadron reconstruction and PV association. In the graph, solid lines (true edges) are shown with their class labels and indicate a physical relationship between nodes: for PV-track edges, the track is associated with the corresponding PV, while for track-track edges, the tracks originate from the same beauty-hadron decay. Dashed lines (false edges) denote examples of connections where no such relationship exists and have a label 0. (b) Heterogeneous modification in blue to the GNN layer updates from Battaglia et al[45] in red. Here, the inputs to the HGNN layer are the sets of node features for tracks and PVs, $V_{\rm tr}$ and $V_{\rm pv}$, sets of edge features for track-track and PV-track edges, $E_{\rm tr}$ and $E_{\rm pv-tr}$, and finally, global features, $u$. The sequence of computations corresponding to Equation 2 is illustrated, while the added pruning tasks are indicated in light brown.
@@
$split e^{\prime k}_{\rm tr}=\phi^{e_{\rm tr}}(e^{k}_{\rm tr},v^{r_{k}}_{\rm tr},v^{s_{k}}_{\rm tr},u)\qquad&\bar{e}^{\prime i}_{\rm tr}=\rho^{e_{\rm tr}\rightarrow v_{\rm tr}}(\{E^{\prime i}_{\rm tr}\})\\ e^{\prime l}_{\rm pv-tr}=\phi^{e_{\rm pv-tr}}(e^{l}_{\rm pv-tr},v^{m_{l}}_{\rm tr},v^{n_{l}}_{\rm pv},u)\qquad&\bar{e}^{\prime i}_{\rm pv-tr}=\rho^{e_{\rm pv-tr}\rightarrow v_{\rm tr}}(\{E^{\prime i}_{\rm pv-tr}\})\\ v^{\prime i}_{\rm tr}=\phi^{v_{\rm tr}}(v^{i}_{\rm tr},\bar{e}^{\prime i}_{\rm tr},\bar{e}^{\prime i}_{\rm pv-tr},u)\qquad&\bar{e}^{\prime j}_{\rm pv-tr}=\rho^{e_{\rm pv-tr}\rightarrow v_{\rm pv}}(\{E^{\prime j}_{\rm pv-tr}\})\\ v^{\prime j}_{\rm pv}=\phi^{v_{\rm pv}}(v^{j}_{\rm pv},\bar{e}^{\prime j}_{\rm pv-tr},u)\qquad&\bar{e}^{\prime}_{\rm tr}=\rho^{e_{\rm tr}\rightarrow u}(\{E^{\prime}_{\rm tr}\}),\bar{v}^{\prime}_{\rm tr}=\rho^{v_{\rm tr}\rightarrow u}(\{V^{\prime}_{\rm tr}\})\\ u^{\prime}=\phi^{u}(\bar{e}^{\prime}_{\rm tr},\bar{e}^{\prime}_{\rm pv-tr},\bar{v}^{\prime}_{\rm tr},\bar{v}^{\prime}_{\rm pv},u)\qquad&\bar{e}^{\prime}_{\rm pv-tr}=\rho^{e_{\rm pv-tr}\rightarrow u}(\{E^{\prime}_{\rm pv-tr}\}),\bar{v}^{\prime}_{\rm pv}=\rho^{v_{\rm pv}\rightarrow u}(\{V^{\prime}_{\rm pv}\})\;,$ (2)
@@
Pruning can be performed either via a top-$k$ operation or by introducing a threshold $y_{\rm cut}$. In this paper, we adopt the latter approach, given its better observed performance and greater interpretability as the threshold $y_{\rm cut}$ directly defines the purity of the retained subgraph, i.e. the fraction of nodes and edges considered relevant according to their predicted scores. During training, these scores are used as continuous weights within the HGNN aggregation functions rather than for hard removal, effectively encouraging the model to favor sparse and informative connections while retaining differentiability. The resulting sparsity bias leads to improved reconstruction performance by suppressing redundant background contributions. At inference time, a hard pruning with $y_{\rm cut}$ is applied to remove redundant background edges and nodes while further increasing processing speed and reduce memory usage. We acknowledge that integrating pruning into training (i.e. pruning-aware or dynamic graph construction) could further improve reconstruction quality. Nonetheless, our empirical results show that incorporating additional MLPs per node and edge already improves the algorithm’s performance, as detailed in the subsequent sections.
@@ 3.3 HGNN architecture and objective loss
$split\mathcal{L}=&\mathcal{L}_{\rm CE}\left(e_{\rm out},y^{\rm LCA}\right)+\beta^{e_{\rm tr}}\sum^{n^{\rm GNN}_{\rm layers}}_{i}\mathcal{L}_{\rm BCE}\left(\hat{y}^{e_{\rm tr}}_{i},y^{e_{\rm tr}}_{i}\right)+\beta^{v_{\rm tr}}\sum^{n^{\rm GNN}_{\rm layers}}_{i}\mathcal{L}_{\rm BCE}\left(\hat{y}^{v_{\rm tr}}_{i},y^{v_{\rm tr}}_{i}\right)\\ &+\beta^{e_{\rm pv-tr}}\sum^{n^{\rm GNN}_{\rm layers}}_{i}\mathcal{L}_{\rm BCE}\left(\hat{y}^{e_{\rm pv-tr}}_{i},y^{e_{\rm pv-tr}}_{i}\right)$ (3)
@@ 3.4 Datasets and training procedure
As depicted in Figure 1, collision events are represented as a heterogeneous graph. Each node and edge type has an input-feature representation. For charged particles (tracks), the input features include the track origin point, $(x_{\rm tr},y_{\rm tr},z_{\rm tr})$, track momenta ($p_{x},p_{y},p_{z}$) and track charge $q$. Meanwhile, PV nodes are represented by their position coordinates, $(x_{\rm pv},y_{\rm pv},z_{\rm pv})$. In contrast, for the GNN benchmarks, we adopted DFEI’s approach of appending the coordinates of the PV with the minimum impact parameter to the features of each track. Although highly discriminating timing information is anticipated for PV association in LHCb Upgrade II [16], this study limited its scope to using only the previously defined positional and kinematic features.
@@
A fully connected graph is defined between all track and PV nodes with track-track and PV-track edges. The input edge representation for PV-track edges includes only the impact parameter of the track with respect to the PV. For track-track edge features include the angle $\theta$ between the three-momentum directions of the two particles, the difference $\Delta z_{\rm tr}$ in origin $z_{\rm tr}$, a Boolean indicating a shared PV according to the minimum impact parameter, and the momentum-transverse distance, which is the distance between the origin points in a plane transverse to the momentum direction. We adopted a loose prefiltering of track-track edges with 99% efficiency from the DFEI [22], which requires edges to satisfy $\theta<0.26$ rad. or the condition of a shared PV according to the minimum impact parameter. For the remaining edges, the ground-truth LCAG is determined $y^{\rm LCA}$, where $y^{\rm LCA}=0$ indicates no shared ancestors. The non-zero values were restricted to $y^{\rm LCA}=1,2,3$, which were found to be sufficient to represent the decay hierarchies present in the simulation dataset. Finally, global input features include the multiplicity of tracks and PVs in the event.
@@
The training and validation datasets consisted of 40,000 and 10,000 events respectively, which are equivalent to those used in our previous work. Nominal training for the ablation studies was performed on an NVIDIA L40S GPU with a batch size of 12. For ablation studies, the base schedule consisted of 30 epochs at a learning rate of $\alpha=10^{-3}$, followed by two epochs at $\alpha=10^{-4}$. Before evaluating the reconstruction performance of the selected architectures, we performed an additional fine-tuning phase on the same dataset, consisting of two epochs with a learning rate of $\alpha=5\times 10^{-5}$ followed by two epochs with $\alpha=2\times 10^{-5}$. During all training phases, training and validation losses as well as classification accuracies were monitored (see  A for examples of training and validation loss curves). The final performance was measured using a test set of 10,000 independent inclusive beauty-hadron decay events and multiple exclusive decay datasets, each containing 5,000 events.
@@ 4.1 Ablation studies
[…]
@@
Figure 2: (a) and (b) Receiver Operating Characteristic curves for edge and node scores $\hat{y}^{e_{\rm tr}}$ and $\hat{y}^{v_{\rm tr}}$ at layers 1, 2, 3 and 8. The pruning selections used in the subsequent results are highlighted. (c) Confusion matrix (in percent) between the number of particles originating from beauty hadrons and the number of particles selected with a tight edge and node pruning, $\hat{y}^{e_{\rm tr}}_{8}>0.2$ and $\hat{y}^{v_{\rm tr}}_{8}>0.2$. (d) Track signal efficiency and background retention as functions of track multiplicity for tight and loose ($\hat{y}^{e_{\rm tr}}_{8}>0.01$ and $\hat{y}^{v_{\rm tr}}_{8}>0.01$) pruning selections.
4.2 Pruning performance
Figure 2(a) and (b) show the discrimination power of the track edge and node scores $\hat{y}^{e_{\rm tr}}$ and $\hat{y}^{v_{\rm tr}}$ at HGNN layers 1, 2, 3 and 8. Here, signal edges or nodes (tracks) refer to those originating from a beauty-hadron decay, whereas background edges and nodes correspond to those not associated with such decays. The discrimination power increases significantly between layers 1 and 2 particularly for the edges, with the area under the curve (AUC) increasing from 0.973 to 0.999. While the discrimination power saturates with subsequent layers, there are marginal increases in the discrimination power for each subsequent HGNN layer.
Figure 2(c) and (d) further quantify the performance of the pruning procedure after applying a selection requiring the LCAG class prediction to be above zero. In particular, Figure 2(c) effectively serves as a confusion matrix, expressed in percent, illustrating how the number of true tracks from beauty hadrons compares with the number of tracks selected under the tight pruning requirement, $\hat{y}^{e_{\rm tr}}_{8}>0.2$ and $\hat{y}^{v_{\rm tr}}_{8}>0.2$. The diagonal dominance demonstrates that the pruning selection preserves the correct track multiplicity for most beauty decays while effectively suppressing spurious tracks. Small off-diagonal entries reflect cases where either additional non-beauty tracks are incorrectly selected or some beauty tracks are missed. Meanwhile, Figure 2(d) compares the efficiencies for signal and background tracks as a function of track multiplicity, using the tight pruning selection and a looser pruning selection with $\hat{y}^{e_{\rm tr}}_{8}>0.01$ and $\hat{y}^{v_{\rm tr}}_{8}>0.01$. The efficiencies are defined as the ratio of surviving signal or background tracks after pruning to those before the pruning selection. Tight pruning strongly suppresses background tracks, particularly at high multiplicities, while retaining a large and stable signal efficiency with increasing event complexity. The looser selection preserves nearly full signal efficiency but admits substantially more background, illustrating the trade-off between purity and efficiency achieved by the pruning criteria.
Figure 3: GNN and HGNN GPU (left) and CPU (right) mean inference time as a function of track multiplicity per event with and without early pruning ($\hat{y}^{e_{\rm tr}/v_{\rm tr}}_{1-3}>10^{-4}$).The error bands indicate $\pm 1$ standard deviation of the inference time distribution for events in each bin.
[…]